THE ROVIBRATIONAL SPECTRUM OF PYRIDINE BETWEEN 600-1300 cm^{−1}

\underline{SIEGHARD ALBERT}, KAREN KEPPLER ALBERT, MARTIN QUACK, PHYSICAL CHEM-ISTRY, ETH ZÜRICH, CH-8093 ZÜRICH, SWITZERLAND; RYAN P.A. BETTENS, DEPART-MENT OF CHEMISTRY, NATIONAL UNIVERSITY OF SINGAPORE, SINGAPORE 117543, SINGAPORE; FRANK C. DE LUCIA, DEPARTMENT OF PHYSICS, THE OHIO STATE UNI-VERSITY, COLUMBUS, OHIO 43210, U.S.A..

One of the great challenges of modern high resolution infrared spectroscopy is the recording and the analysis of rotationally resolved spectra of relatively heavy molecules containing numerous vibrational modes. Here we present new results on the prototype heterocyclic molecule of pyridine. This molecule is of astrophysical interest and can be related to other nitrogen containing biological aromatic systems. Up to now only vibrationally resolved infrared spectra and rotational spectra exist, including recent FASSST spectra^{a}. We have recorded the spectrum of pyridine in the region 600-1300 cm^{−1} with our Zürich Bruker 2001 prototype Fourier transform infrared spectrometer^{b} with an instrumental resolution of 0.0007 cm^{−1}, essentially Doppler limited, with Doppler widths ranging from 0.0008 to 0.0018 cm^{−1}. The analysis of the bands located at ν̃_{c} = 700.2528 cm^{−1} (ν_{11}), ν̃_{c} = 744.0046 cm^{−1} (ν_{4}) and ν̃_{c} = 1031.6339 cm^{−1} (ν_{12}) will be discussed. Perturbations of the ν_{1} and ν_{18a} modes have been observed. We will give an outlook how in the future the rovibrational spectra of larger molecules of biological relevance can be recorded and analysed using the combination of our highly resolving spectrometer with Synchrotron radiation and applying our spectral analysis tools to the complicated spectra.

aE. Ye, R.P.A. Bettens, F.C. De Lucia, D.T. Petkie and S. Albert, J. Mol. Spectrosc., submitted.

bS. Albert, K.K. Albert, and M. Quack, Trends in Optics and Photonics, 84 177 (2003), S. Albert, H. Hollenstein, M. Quack and M. Willeke, Mol. Phys., 102, 1671 (2004).